OpenStudy (anonymous):

Solve the following inequalities: a. x^2−x−6≤0 b. x^3+5x^2−4x−20>0

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

Where are you getting all these problems? They are so wide ranging in content.

OpenStudy (mertsj):

Factor. And remember if you multiply two factors and the product is less than 0, then 1 factor must be negative and one must be positive.

OpenStudy (anonymous):

can you show me an example of what you mean ? @Mertsj

OpenStudy (mertsj):

If xy is negative then either x is negative and y positive or x is positive and y is negative.

OpenStudy (anonymous):

ok then can you do a pleaseeeeeeee @Mertsj

OpenStudy (anonymous):

then I'll know what your talking about

OpenStudy (mertsj):

(x-3)(x+2) is negative or 0 so x-3<0 and x+2>0 or x-3>0 and x+2<0 The first case: x<3 and x>-2 Second case: x>3 and x<-2 Obviously that is impossible since a number cannot be both greater than 3 and less than -2 so the answer is the first case.

OpenStudy (anonymous):

I can't factor b @Mertsj

OpenStudy (mertsj):

Factor x^2 out of the first 2 terms Factor -4 out of the last 2 terms.

OpenStudy (anonymous):

okay I did that @Mertsj

OpenStudy (anonymous):

x^2(x+5)-4(x-5)>0

OpenStudy (anonymous):

now what???

OpenStudy (mertsj):

-4x-20=-4(x+5)

OpenStudy (anonymous):

huuh

OpenStudy (mertsj):

Where did you get -4(x-5) ???

OpenStudy (mertsj):

x^2(x+5)-4(x-5)>0 Where did you get that?

OpenStudy (anonymous):

factored out -4

OpenStudy (mertsj):

Exactly my point. You did that incorrectly.

OpenStudy (anonymous):

you said Factor -4 out of the last 2 terms

OpenStudy (mertsj):

As I said: -4x-20=-4(x+5)

OpenStudy (anonymous):

oh ok I put - instead +

OpenStudy (mertsj):

So now your two terms have a common factor of x+5. Factor it out.

OpenStudy (anonymous):

how did you get -4x-20?

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

Perhaps you would care to read the original problem?

OpenStudy (anonymous):

I factored it now what?

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

The same as before. If you multiply two numbers and the answer is positive then both factors must be negative or both factors must be positive.

OpenStudy (anonymous):

Im having trouble

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

Write what you did so far.

OpenStudy (anonymous):

x+5 >0

OpenStudy (mertsj):

If you multiply two factors and the answer is positive then BOTH factors must be negative or BOTH factors must be positive.

OpenStudy (anonymous):

so the two factors were (x+5)(x-5) ?? right

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

Post your factored form of the polynomial.

OpenStudy (anonymous):

I have three more questions like this, so can you just show this one to me and I'll do the other 3 on my own

OpenStudy (mertsj):

Post your factored form of the polynomial.

OpenStudy (anonymous):

-4(x+5)=-4(x+5)

OpenStudy (mertsj):

You need to refactor the polynomial. The right side is 0. Keep everything else on the left.

OpenStudy (anonymous):

-4(x+5)=0

OpenStudy (mertsj):

If you multiply that out, it does not multiply back to the left side. What ever you write should be the same as x^3+5x^2-4x-20

OpenStudy (anonymous):

(-4x-20)= -4(x+5)

OpenStudy (anonymous):

idk !!!

OpenStudy (mertsj):

What test are you working on?

OpenStudy (anonymous):

huuh?

OpenStudy (anonymous):

what test

OpenStudy (anonymous):

??

OpenStudy (mertsj):

Where are you getting all these problems that are not related?

OpenStudy (anonymous):

they are review questions

OpenStudy (mertsj):

For what course?

OpenStudy (anonymous):

math

OpenStudy (mertsj):

What math?

OpenStudy (anonymous):

polynomial functions unit review

OpenStudy (anonymous):

why????

OpenStudy (mertsj):

x^2(x+5)-4(x-5)>0

OpenStudy (mertsj):

You posted that earlier. I pointed out that the -5 should be +5. Did you ever correct that?

OpenStudy (anonymous):

oh yes.

OpenStudy (mertsj):

So post the corrected version.